OpenStudy (anonymous):

PLEASE HELP What is the 41st term of the arithmetic sequence where a1 = 15 and a15 = –41 ?

OpenStudy (diyadiya):

\[\large a_n= a_1+ (n-1)d\]\[\large a_1=15\]\[\large a_{15}=a_1+14d\]

OpenStudy (diyadiya):

\[\large a_{15}=-41\]\[\large a_1=15\] Substitute and find d

OpenStudy (diyadiya):

\[\large \text{After finding d ,We know that}\]\[\large a_{41}=a+(41-1)d=a+40d\] Substitute a1and d & find a41

OpenStudy (anonymous):

wait so do i solve it like that or? im sorry Im confused